OpenStudy (anonymous):

How would you solve this? A car accelerates (from rest) at a rate of 4.5 m/s2 for 3 s. How far did it travel in this time? Select one: a. 13.5 m b. 20.25 m c. 7.5 m d. 4.5 m

OpenStudy (anonymous):

*4.5m/s^2

OpenStudy (anonymous):

X= Vot+1/2a(t^2) X= 1/2(4.5m/s^2)(3^2) X= 20.25m

OpenStudy (anonymous):

Awesome thanks so much dude ! :)

OpenStudy (anonymous):

You're welcome :3